OpenStudy (lacedupcountry):

I need someone to help me with my homework, Just to check if I'm doing it right. I'll give medals!

OpenStudy (michele_laino):

please your starting prices are: A. =25, B=22, C=30 so I think you have pick a starting price less than 25, are you agree?

OpenStudy (lacedupcountry):

Ohh okay, yeah i agree i totally messed that up @Michele_Laino sorry i didn't get back sooner i was taking another test so i could do 23?

OpenStudy (lacedupcountry):

So i changed question 1: F(x)=23(1.3)x

OpenStudy (swagster23780):

I agree with Michele_Laino

OpenStudy (lacedupcountry):

Yes and i did that, I got rid of all of my other answers and made a new Function. Now can someone help me out the rest of the way please? With number 2?

OpenStudy (lacedupcountry):

2. Graph the price function for Stock D. Explain the key features of the graph with complete sentences. so would i write it like this: F(1)=23(1.3)1 F(1)=29.9 F(2)=23(1.3)2 F(2)=59.8 F(x)=59.8x

OpenStudy (mathmath333):

for \(\large\tt \color{red}{A}\) \[\large f(x)=25(1.08)^x\] \(\large\tt \color{black}{A~~initial ~~value =25}\) it is equal to \(\Large\tt \color{blacck}{increase}\) in for \(\Large\tt \color{blacck}{8\%}\) per day for \(\large\tt \color{red}{B}\) \(\large\tt \color{black}{B~~initial ~~value =22}\) \(\large\tt \color{black}{21.56=22(0.98) =22(1-0.02)}\) hence for B it will be \(\Large\tt \color{blacck}{decrease}\) in \(\Large\tt \color{blacck}{2\%}\) per day for \(\large\tt \color{red}{C}\) it is already given \(\large\tt \color{black}{C~~initial ~~value =30}\) it is equal to \(\Large\tt \color{blacck}{increase}\) in for \(\Large\tt \color{blacck}{4\%}\) per day for\(\large\tt \color{red}{D}\) second lowest starting price=\(\large\tt \color{black}{25}\) second highest rate of growth=\(\large\tt \color{black}{4\%}\) there for function for D will be \(\Large\tt \color{black}{f(x)=25(1.04)^x}\) x f(x) 0 24 1 26 2 27.04

OpenStudy (lacedupcountry):

Oh my gosh thank you so much, this makes a lot more sense now!

OpenStudy (mathmath333):

i solved this type of question months ago see this for graph http://openstudy.com/users/aussie246#/updates/543ae001e4b01d3397661494

OpenStudy (lacedupcountry):

Okay thankyou!